Four identical metallic objects carry the following charges: +1.45, +6.97, -4.56, and -9.52C. The objects are brought simultaneously into contact, so that each touches the others. Then they are separated. (a) What is the final charge on each object? (b) How many electrons (or protons) make up the final charge on each object?
Answer #1

(a)
the first answer is correct, total the charges and divide by how many there are
+1.45C+6.97C+(-4.56C)+(-9.52C) = -5. 66C
-5. 66C/4 = -1.415C

(b)
my problem is given in µC which is = _ x 10^-6 C
if it's C though just divide the total charge of the object by the charge of one electron (1.6 x 10^-19 C)
-1.415C/1.6x10^-19C = -8. 8x10^18C
the negative sign in front tells you that the final charge is made up of electrons